OpenStudy (rock_mit182):

For each pair of functions f and g below, find f(g(x)) and g(f(x)) . Then, determine whether f and g are inverses of each other. f(x) = \[\frac{ 2}{ x } x \neq 2\] g(x)= \[\frac{ 2}{ x } x \neq 2\] f(g(x)) = g(f(x)) =

OpenStudy (rock_mit182):

f(g(x)) = 2/(2)/x = x g(f(x)) = 2/(2)/(x) = x They are inverses of each other

OpenStudy (rock_mit182):

is that right ?